# Beale–Kato–Majda-type continuation criteria for Hall- and electron-magnetohydrodynamics

### v1

@@ 1. Introduction
$\nabla\cdot{\bf B}$ $=0,$
where ${\bf B}(t):\mathbb{R}^{3}_{x}\to\mathbb{R}^{3}$ for each $t\geq 0$. By vector calculus identities, the first equation can be rewritten as
$(\partial_{t}-\Delta){\bf B}+{\bf B}\cdot\nabla(\nabla\times{\bf B})-(\nabla\times{\bf B})\cdot\nabla{\bf B}=0.$
The Hall MHD equation (H-MHD) with resistivity takes the form
@@
$\nabla\cdot{\bf u}=\nabla\cdot{\bf B}$ $=0$
where $\nu\geq 0$ and ${\bf u}(t),{\bf B}(t):\mathbb{R}^{3}_{x}\to\mathbb{R}^{3}$ for each $t\geq 0$. By vector calculus identities, the equation for $\partial_{t}{\bf B}$ can be rewritten as
@@
(1.3) $\int_{0}^{T}\|{\nabla{\bf J}=(t)}\|_{L^{\infty}}\,\mathrm{d}t<+\infty,$
@@
2. Mathematically, a novel aspect of these criteria is that (1.3) and (1.4) impose very little control on the low frequency part of ${\bf B}$ due to the presence of several derivatives (i.e., $\nabla{\bf J}=\nabla\nabla\times{\bf B}$); see also the brief review of the previous results below. Indeed, obtaining a sufficiently strong control of ${\bf B}$ itself starting from (1.3) and (1.4) is the main difficulty of the proofs. We address this by making a (seemingly new) observation that (1.3) and (1.4) leads to a parabolic maximum principle for $|{{\bf B}}|^{2}$, and hence controls the growth of $\|{{\bf B}(t)}\|_{L^{\infty}}$. See Step 1 in Sections 3 and 4.
3. Since ×$\nabla\times{\bf J}$ is divergence-free, it is reasonable to suspect whether (×$\nabla(\nabla\times{\bf J})$ in Theorem 1.1 can be replaced by ×(×$\nabla\times(\nabla\times{\bf J})$. Unfortunately, our proof breaks down, specifically the maximum principle for $|{{\bf B}}|^{2}$. Same comment applies to the question of generalizing Theorem 1.2 by replacing $\nabla({\bf u}-{\bf J})$ by $\boldsymbol{\omega}-\nabla\times{\bf J}$.
4. As remarked earlier, (1.3) is also invariant under the scalings of the non-resistive E-MHD, i.e., (1.1) without $\Delta{\bf B}$ on the right-hand side. One may therefore ask whether Theorem 1.1 holds even in the non-resistive case. The answer is likely negative in view of the recent illposedness results of Jeong and the second author [11], in the following sense. As is usual, our proof of Theorem 1.1 proceeds by establishing persistence of regularity, i.e., showing that a sufficiently high Sobolev norm of ${\bf B}$ (concretely, the $H^{4}$-norm) enjoys a uniform bound for $0\leq t<T$ in terms of the control norm in (1.3). Since the time integral of the $H^{4}$-norm bounds the control norm in (1.3), such a persistence of regularity result would lead to an short time a-priori $H^{4}$-norm bound. However, this is impossible in view of the illposedness results in [11]. A similar remark applies to Hall–MHD.
@@
We end the introduction with a brief discussion of previously known continuation criteria for (1.1) and (1.2). The local well-posednes of classical solution of (1.2) (and (1.1)) is known thanks to the work [3] of Chae, Degond, and Liu. Moreover, the authors established the following continuation criterion for the local smooth solution beyond the time $T$,
@@
with
$\frac{3}{p}+\frac{2}{q}\leq 1,\qquad\frac{3}{\beta}+\frac{2}{\gamma}\leq 1\qquad\mbox{ and }\qquad p,\beta\in(3,\infty].$
In the limit case $q=\gamma=2$, the authors also showed that if
@@
The first author established a regularity criterion for the 3D Hall-MHD (1.2) with $\nu=1$ on $\mathbb{T}^{3}$ in term of low frequencies of $\nabla{\bf u}$ and $\nabla^{2}{\bf B}$ in [5]. The author defined time dependent dissipation wavenumber $\kappa_{{\bf u}}(t)$ and $\kappa_{{\bf B}}(t)$ with respect to ${\bf u}$ and ${\bf B}$. Assuming
[…]
the local smooth solution $({\bf u}(t),{\bf B}(t))$ can be extended over $T$. The projection notation $P_{<}$ is introduced in Section 2. It was shown in [5] that the low frequency condition (1.8) is weaker than (1.6) and (1.7).We also mention the articles [8, 13], in which a continuation criterion for (1.2) with [^2] We note that these proofs rely heavily on the fact that the variable ${\bf Z}:={\bf B}+\boldsymbol{\omega}$ solves an advection–diffusion equation (with the advection given by ${\bf u}$), which holds only when $\nu=1$ (i.e., viscosity and resistivity match). $\nu=1$ only in terms of the ${\bf u}$ was derived.
@@ 2. Preliminaries
We employ the following convention for Littlewood–Paley projections. Fix $m_{<0}\in C^{\infty}_{c}(\mathbb{R}^{3})$ that is radial, non-increasing (in $r$), $\operatorname{supp}m_{<0}\subseteq\{\xi\in\mathbb{R}^{3}:|{\xi}|<2\}$ and $m_{<0}(\xi)=1$ on $\{\xi\in\mathbb{R}^{3}:|{\xi}|<1\}$. We define $P_{<k}f:=\mathcal{F}^{-1}[m_{<0}(2^{-k}\xi)\mathcal{F}[f]]$, and $P_{k}f:=P_{<k+1}f-P_{<k}f$ (where $\mathcal{F}$ is the Fourier transform).
@@
where each $c_{I^{\prime},I^{\prime\prime}}$ is some combinatorial coefficient. Taking the inner product with $\partial^{I}{\bf B}$, integrating on $\mathbb{R}^{3}$ and summing over all $I$ such that $|{I}|=4$, we obtain
@@
$\lesssim\eta\|{\nabla^{5}{\bf B}}\|_{L^{2}}^{2}+\left[\left(\log(2+\|{\nabla^{4}{\bf B}}\|_{L^{2}})+(1+\eta^{-1})\|{{\bf B}}\|_{L^{\infty}}\right)\|{\nabla{\bf J}}\|_{L^{\infty}}+\|{{\bf B}}\|_{L^{\infty}}\right]\|{\nabla^{4}{\bf B}}\|_{L^{2}}^{2}$
@@
Indeed, taking the dot product of the second equation of (1.2) with $B$ yields
@@
$\begin{split}&\frac{\mathrm{d}}{\mathrm{d}t}\left(\|{\nabla^{4}{\bf u}}\|_{L^{2}}^{2}+\|{\nabla^{4}{\bf B}}\|_{L^{2}}^{2}\right)+\nu\|{\nabla^{5}{\bf u}}\|_{L^{2}}^{2}+\|{\nabla^{5}{\bf B}}\|_{L^{2}}^{2}\\ \lesssim&\left(\log(2+\|{\nabla^{4}{\bf u}}\|_{L^{2}}+\|{\nabla^{4}{\bf B}}\|_{L^{2}})+(1+\eta^{-1})\|{{\bf B}}\|_{L^{\infty}}+\|{{\bf u}}\|_{L^{2}}+1\right)\\ &\cdot\left(\|{\boldsymbol{\omega}}\|_{L^{\infty}}+\|{\nabla(u-{\bf J})}\|_{L^{\infty}}+\|{{\bf B}}\|_{L^{\infty}}\right)\left(\|{\nabla^{4}{\bf u}}\|_{L^{2}}^{2}+\|{\nabla^{4}{\bf B}}\|_{L^{2}}^{2}\right)\\ \end{split}$
In view of the condition (1.4), it then follows from the generalized Grönwall’s inequality that $\|{\nabla^{4}{\bf u}(t)}\|_{L^{2}}^{2}+\|{\nabla^{4}{\bf B}(t)}\|_{L^{2}}^{2}$ is uniformly bounded up to the time $t=T$. ∎

### v2

@@ 1. Introduction
$\nabla\cdot{\bf B}$ $=0$
where ${\bf B}(t):\mathbb{R}^{3}_{x}\to\mathbb{R}^{3}$ for each $t\geq 0$. The Hall MHD equation (H-MHD) with resistivity takes the form
@@
$\nabla\cdot{\bf u}$ $=0,$
$\nabla\cdot{\bf B}$ $=0$
where $\nu\geq 0$ is the constant viscosity parameter and unknowns ${\bf u}(t),{\bf B}(t):\mathbb{R}^{3}_{x}\to\mathbb{R}^{3}$, $p(t):\mathbb{R}^{3}_{x}\to\mathbb{R}$ for each $t\geq 0$. Both systems (1.1) and (1.2) are models from plasma physics, describing respectively the evolution of magnetic field ${\bf B}$ without or with the influence of the velocity field ${\bf u}$ of the ion flow. System (1.1) is regarded as an approximating model of (1.2) when ${\bf u}$ is negligible and the constraint $-{\bf B}\cdot\nabla{\bf B}+\nabla p=0$ is relaxed. By vector calculus identities, the first equation of (1.1) can be rewritten as
$(\partial_{t}-\Delta){\bf B}+{\bf B}\cdot\nabla(\nabla\times{\bf B})-(\nabla\times{\bf B})\cdot\nabla{\bf B}=0,$
while the equation for $\partial_{t}{\bf B}$ in (1.2) can be rewritten as
@@
(1.3) $\int_{0}^{T}\|{\nabla{\bf J}(t)}\|_{L^{\infty}}\,\mathrm{d}t<+\infty,$
@@
2. Mathematically, a novel aspect of these criteria is that (1.3) and (1.4) impose very little control on the low frequency part of ${\bf B}$ due to the presence of several derivatives (i.e., $\nabla{\bf J}=\nabla\nabla\times{\bf B}$); see also the brief review of the previous results below. Indeed, obtaining a sufficiently strong control of ${\bf B}$ itself starting from (1.3) and (1.4) is the main difficulty of the proofs. We address this by making a (seemingly new) observation that (1.3) or (1.4) leads to a parabolic maximum principle for $|{{\bf B}}|^{2}$, and hence controls the growth of $\|{{\bf B}(t)}\|_{L^{\infty}}$. See Step 1 in Sections 3 and 4.
3. Since ${\bf J}$ is divergence-free, it is reasonable to suspect whether $\nabla{\bf J}$ in Theorem 1.1 can be replaced by ×$\nabla\times{\bf J}$. Unfortunately, our proof breaks down, specifically the maximum principle for $|{{\bf B}}|^{2}$. Same comment applies to the question of generalizing Theorem 1.2 by replacing $\nabla({\bf u}-{\bf J})$ with $\boldsymbol{\omega}-\nabla\times{\bf J}$.
4. As remarked earlier, (1.3) is also invariant under the scalings of the non-resistive E-MHD, i.e., (1.1) without $\Delta{\bf B}$ on the right-hand side. One may therefore ask whether Theorem 1.1 holds even in the non-resistive case. The answer is likely negative in view of the recent illposedness results of Jeong and the second author [11], in the following sense. As is usual, our proof of Theorem 1.1 proceeds by establishing persistence of regularity, i.e., showing that a sufficiently high Sobolev norm of ${\bf B}$ (concretely, the $H^{4}$-norm) enjoys a uniform bound for $0\leq t<T$ in terms of the control norm in (1.3). Since the time integral of the $H^{4}$-norm bounds the control norm in (1.3), such a persistence of regularity result would lead to a short time a-priori $H^{4}$-norm bound. However, this is impossible in view of the illposedness results in [11]. A similar remark applies to Hall–MHD.
@@
We end the introduction with a brief discussion of previously known continuation criteria for (1.1) and (1.2). The local well-posedness of classical solution of (1.2) (and (1.1)) is known thanks to the work [3] of Chae, Degond, and Liu. Moreover, the authors established the following continuation criterion for the local smooth solution beyond the time $T$,
@@
with $\frac{3}{p}+\frac{2}{q}\leq 1$, $\frac{3}{\beta}+\frac{2}{\gamma}\leq 1$ and $p,\beta\in(3,\infty]$. In the limit case $q=\gamma=2$, the authors also showed that if
@@
The first author established a regularity criterion for the 3D Hall-MHD (1.2) with $\nu=1$ on $\mathbb{T}^{3}$ in term of low frequencies of $\nabla{\bf u}$ and $\nabla{\bf B}$ in [5]. The author defined time dependent dissipation wavenumber $\kappa_{{\bf u}}(t)$ and $\kappa_{{\bf B}}(t)$ with respect to ${\bf u}$ and ${\bf B}$. Assuming
(1.8) $\int_{0}^{T}\left(\|P_{<\kappa_{{\bf u}}(t)}\nabla{\bf u}(t)\|_{B^{0,\infty}_{\infty}}+2^{\kappa_{{\bf B}}(t)}\|P_{<\kappa_{{\bf B}}(t)}\nabla{\bf B}(t)\|_{B^{0,\infty}_{\infty}}\right)\,\mathrm{d}t<\infty,$
the local smooth solution $({\bf u}(t),{\bf B}(t))$ can be extended over $T$. The projection notation $P_{<}$ is introduced in Section 2. It was shown in [5] that the low frequency condition (1.8) is weaker than (1.6) and (1.7); nevertheless, the new criterion (1.4) does not imply (1.8), and vice versa. We also mention the articles [8, 13], in which a continuation criterion for (1.2) with [^2] We note that these proofs rely heavily on the fact that the variable ${\bf Z}:={\bf B}+\boldsymbol{\omega}$ solves an advection–diffusion equation (with the advection given by ${\bf u}$), which holds only when $\nu=1$ (i.e., viscosity and resistivity match). $\nu=1$ only in terms of the ${\bf u}$ was derived.
@@ 2. Preliminaries
We employ the following convention for Littlewood–Paley projections. Fix the function $m_{<0}\in C^{\infty}_{c}(\mathbb{R}^{3})$ that is radial and non-increasing (in $r$) with $\operatorname{supp}m_{<0}\subseteq\{\xi\in\mathbb{R}^{3}:|{\xi}|<2\}$ and $m_{<0}(\xi)=1$ on $\{\xi\in\mathbb{R}^{3}:|{\xi}|<1\}$. We define $P_{<k}f:=\mathcal{F}^{-1}[m_{<0}(2^{-k}\xi)\mathcal{F}[f]]$, and $P_{k}f:=P_{<k+1}f-P_{<k}f$ (where $\mathcal{F}$ is the Fourier transform).
@@
where each $c_{I^{\prime},I^{\prime\prime}}$ is some combinatorial coefficient. Taking the inner product of the equation above with $\partial^{I}{\bf B}$, integrating on $\mathbb{R}^{3}$ and summing over all $I$ with $|{I}|=4$, we obtain
@@
$\lesssim\eta\|{\nabla^{5}{\bf B}}\|_{L^{2}}^{2}+\left[\left(\log(2+\|{\nabla^{4}{\bf B}}\|_{L^{2}})+(1+\eta^{-1})\|{{\bf B}}\|_{L^{\infty}}\right)\|{\nabla{\bf J}}\|_{L^{\infty}}+\|{{\bf B}}\|_{L^{\infty}}\right]\|{\nabla^{4}{\bf B}}\|_{L^{2}}^{2}.$
@@
Indeed, taking the dot product of the second equation of (1.2) with ${\bf B}$ yields
@@
$\begin{split}&\frac{\mathrm{d}}{\mathrm{d}t}\left(\|{\nabla^{4}{\bf u}}\|_{L^{2}}^{2}+\|{\nabla^{4}{\bf B}}\|_{L^{2}}^{2}\right)+\nu\|{\nabla^{5}{\bf u}}\|_{L^{2}}^{2}+\|{\nabla^{5}{\bf B}}\|_{L^{2}}^{2}\\ \lesssim&\left(\log(2+\|{\nabla^{4}{\bf u}}\|_{L^{2}}+\|{\nabla^{4}{\bf B}}\|_{L^{2}})+(1+\eta^{-1})\|{{\bf B}}\|_{L^{\infty}}+\|{{\bf u}}\|_{L^{2}}+1\right)\\ &\cdot\left(\|{\boldsymbol{\omega}}\|_{L^{\infty}}+\|{\nabla(u-{\bf J})}\|_{L^{\infty}}+\|{{\bf B}}\|_{L^{\infty}}\right)\left(\|{\nabla^{4}{\bf u}}\|_{L^{2}}^{2}+\|{\nabla^{4}{\bf B}}\|_{L^{2}}^{2}\right).\end{split}$
It is noteworthy that we did not use the term $\nu\|{\nabla^{5}{\bf u}}\|_{L^{2}}^{2}$ in the estimates above and hence the inequality holds for all $\nu\geq 0$, including the inviscid case $\nu=0$. In view of the condition (1.4), it then follows from the generalized Grönwall’s inequality that $\|{\nabla^{4}{\bf u}(t)}\|_{L^{2}}^{2}+\|{\nabla^{4}{\bf B}(t)}\|_{L^{2}}^{2}$ is uniformly bounded up to the time $t=T$. ∎